# Transparent Networks for Multivariate Time Series

### v2

@@ Abstract
Transparent models, which are machine learning models that produce inherently interpretable predictions, are receiving significant attention in high-stakes domains. However, despite much real-world data being collected as time series, there is a lack of studies on transparent time series models. To address this gap, we propose a novel transparent neural network model for time series called Generalized Additive Time Series Model (GATSM). GATSM consists of two parts: 1) independent feature networks to learn feature representations, and 2) a transparent temporal module to learn temporal patterns across different time steps using the feature representations. This structure allows GATSM to effectively capture temporal patterns and handle dynamic-length time series while preserving transparency. Empirical experiments show that GATSM significantly outperforms existing generalized additive models and achieves comparable performance to black-box time series models, such as recurrent neural networks and Transformer. In addition, we demonstrate that GATSM finds interesting patterns in time series. The source code is available at https://github.com/gim4855744/GATSM.
@@ 1 Introduction
Artificial neural networks excel at learning complex representations and demonstrate remarkable predictive performance across various fields. However, their complexity makes interpreting the decision-making processes of neural network models challenging. Consequently, post-hoc explainable artificial intelligence (XAI) methods, which explain the predictions of trained black-box models, have been widely studied in recent years [1, 2, 3, 4]. XAI methods are generally effective at providing humans with understandable explanations of model predictions. However, they may produce incorrect and unfaithful explanations of the underlying black-box model and cannot provide actual contributions of input features to model predictions [5, 6]. Therefore, their applicability to high-stakes domains-such as healthcare and fraud detection, where faithfulness to the underlying model and actual contributions of features are important-is limited.
@@
Due to these limitations, transparent (i.e., inherently interpretable) models are attracting attention as alternatives to XAI in high-stakes domains [7, 8, 9]. Modern transparent models typically adhere to the generalized additive model (GAM) framework [10]. A GAM consists of independent functions, each corresponding to an input feature, and makes predictions as a linear combination of these functions (e.g., the sum of all functions). Therefore, each function reflects the contribution of its respective feature. For this reason, interpreting GAMs is straightforward, making them widely used in various fields, such as healthcare [11, 12], survival analysis [13], and model bias discovery [7, 14, 15]. However, despite much real-world data being collected as time series, research on GAMs for time series remains scarce. Consequently, the applicability of GAMs in real-world scenarios is still limited.
@@
To overcome this limitation, we propose a novel transparent model for multivariate time series called Generalized Additive Time Series Model (GATSM). GATSM consists of independent feature networks to learn feature representations and a transparent temporal module to learn temporal patterns. Since employing distinct networks across different time steps requires a massive amount of learnable parameters, the feature networks in GATSM share the weights across all time steps, while the temporal module independently learns temporal patterns. GATSM then generates final predictions by integrating the feature representations with the temporal information from the temporal module. This strategy allows GATSM to effectively capture temporal patterns and handle dynamic-length time series while preserving transparency. Additionally, this approach facilitates the separate extraction of time-independent feature contributions, the importance of individual time steps, and time-dependent feature contributions through the feature functions, temporal module, and final prediction. To demonstrate the effectiveness of GATSM, we conducted empirical experiments on various time series datasets. The experimental results show that GATSM significantly outperforms existing GAMs and achieves comparable performances to black-box time series models, such as recurrent neural networks and Transformer [16]. In addition, we provide visualizations of GATSM’s predictions to demonstrate that GATSM finds interesting patterns in time series.
2 Related Works
Various XAI studies have been conducted over the past decade [7, 8, 9, 17, 18]; however, they are less relevant to the transparent model that is the subject of this study. Therefore, we refer readers to [19, 20] for more detailed information on recent XAI research. In this section, we review existing transparent models closely related to our GATSM and discuss their limitations.
@@ 2 Related Works
The simple linear model is designed to fit the conditional expectation $g\left(\mathbb{E}\left(y\mid\textbf{x}\right)\right)=\sum_{i=1}^{M}x_{i}w_{i}$, where $g(\cdot)$ is a link function, $M$ indicates the number of input features, $y$ is the target value for the given input features $\textbf{x}\in\mathbb{R}^{M}$, and $w_{i}\in\mathbb{R}$ is the learnable weight for $x_{i}$. This model captures only linear relationships between the target $y$ and the inputs x. To address this limitation, GAM [10] extends the simple linear model to the generalized form as follows:
@@
where each $f_{i}(\cdot)$ is a function that models the effect of a single feature, referred as a feature function. Typically, $f_{i}\left(\cdot\right)$ becomes a non-linear function such as a decision tree or neural network to capture non-linear relationships.
Originally, GAMs were fitted via the backfitting algorithm using smooth splines [10, 21]. Later, Yin Lou et al. [22] and Harsha Nori et al. [23] have proposed boosted decision tree-based GAMs, which use boosted decision trees as feature functions. Spline- and tree-based GAMs have less flexibility and scalability. Thus, extending them to transfer or multi-task learning is challenging. To overcome this problem, various neural network-based GAMs have been proposed in recent years. Potts [24] introduced generalized additive neural network, which employs 2-layer neural networks as feature functions. Similarly, Rishabh Agarwal et al. [7] proposed neural additive model (NAM) that employs multi-layer neural networks. To improve the scalability of NAM, Chun-Hao Chang et al. [8] and Filip Radenovic et al. [9] proposed the neural oblivious tree-based GAM and the basis network-based GAM, respectively. Xu et al. [25] introduced a sparse version of NAM using the group LASSO. One disadvantage of GAMs is their limited predictive power, which stems from the fact that they only learn first-order feature interactions-i.e., relationships between the target value and individual features. To address this, various studies have been conducted to enhance the predictive powers of GAMs by incorporating higher-order feature interactions, while still maintaining transparency. GA2M [26] simply takes pairwise features as inputs to learn pairwise interactions. GAMI-Net [27], a neural network-based GAM, consists of networks for main effects (i.e., first-order interactions) and pairwise interactions. To enhance the interpretability of GAMI-Net, the sparsity and heredity constraints are added, and trivial features are pruned in the training process. Sparse interaction additive network [28] is a 3-phase method for exploiting higher-order interactions. Initially, a black-box neural network is trained; subsequently, the top-$k$ important features are identified using explainable feature attribution methods like LIME [1] and SHAP [2], and finally, NAM is trained with these extracted features. Dubey et al. [29] introduced scalable polynomial additive model, an end-to-end model that learns higher-order interactions via polynomials. Similarly, Kim et al. [15] proposed higher-order NAM that utilizes the feature crossing technique to capture higher-order interactions. Despite their capabilities, the aforementioned GAMs cannot process time series data, which limits their applicability in real-world scenarios. Recently, neural additive time series model (NATM) [30], a time-series adaptation of NAM, has been proposed. However, NATM handles each time step independently with separate feature networks. This approach cannot capture effective temporal patterns and only takes fixed-length time series as input. Our GATSM not only captures temporal patterns but also handles dynamic-length time series. Table 1 shows the advantages of our GATSM compared to existing GAMs.
Table 1: Advantages of the proposed GATSM.
Time series input Temporal pattern Dynamic time series
existing GAMs
@@
3 Problem Statement
We tackle the problem of the existing GAMs on time series. Equation (1) outlines the GAM framework for tabular data, which fails to capture the interactions between current and previous observations in time series. A straightforward method to extend GAM to time series, adopted in NATM, is applying distinct feature functions to each time step and summing them to produce predictions:
where $\textbf{X}\in\mathbb{R}^{T\times M}$ is a time series with $T$ time steps and $M$ features, and $t$ is the current time step. This method can handle time series data as input but fails to capture effective temporal patterns because the function $f_{i,j}\left(\cdot\right)$ still does not interact with previous time steps. To overcome this problem, we suggest a new form of GAM for time series defined as follows:
Definition 3.1 GAMs for time series, which capture temporal patterns hold the form of Equation 3.
In Equation (3), the function $f\left(\cdot,\cdot\right)$ can capture interactions between current and previous time steps. Therefore, GAMs adhering to Definition 3.1 are capable of capturing temporal patterns. However, implementing such a model while maintaining transparency poses challenges. In the following section, we will describe our approach to implementing a GAM that holds Definition 3.1. To the best of our knowledge, no existing literature addresses Definition 3.1.
4 Our Method: Generalized Additive Time Series Model
Figure 1: Architecture of GATSM.
4.1 Architecture
Figure 1 shows the overall architecture of GATSM. Our model has two modules: 1) feature networks, called time-sharing neural basis model, for learning feature representations, and 2) masked multi-head attention for learning temporal patterns.
Time-Sharing NBM: Assume a time series with $T$ time steps and $M$ features. Applying GAMs to this time series necessitates $T\times M$ feature functions, which becomes problematic when dealing with large $T$ or $M$ due to increased model size. This limits the applicability of GAMs to real-world datasets. To overcome this problem, we extend neural basis model (NBM) [9] to time series as:
@@ 4.1 Architecture
We refer to this extended version of NBM as time-sharing NBM. Time-sharing NBM has $B$ basis functions, with each basis $h_{k}(\cdot)$ taking a feature $x_{i,j}$ as input. The feature-specific weight $w_{j,k}^{nbm}$ then projects the basis to the transformed feature $\tilde{x}_{i,j}$. As depicted in Equation 4, the basis functions are shared across all features and time steps, drastically reducing the number of required feature functions $T\times M$ to $B$. We use $B=100$ and implement $h_{k}\left(\cdot\right)$ using multi-layer perceptron (MLP).
Masked MHA: GATSM employs multi-head attention (MHA) to learn temporal patterns. Although the dot product attention [16] is popular, simple dot operation has low expressive power [31]. Therefore, we adopt the 2-layer attention mechanism proposed by [31] to GATSM. We first transform $\tilde{\textbf{x}}_{i}=\left[\tilde{x}_{i,1},\tilde{x}_{i,2},\cdots,\tilde{x}_{i,M}\right]\in\mathbb{R}^{M}$ produced by Equation 4 as follows:
@@
where $\textbf{Z}\in\mathbb{R}^{M\times D}$ is a learnable weight, $\textbf{pe}_{i}=\left[pe_{i,1},pe_{i,2},\cdots,pe_{i,D}\right]\in\mathbb{R}^{D}$ is the positional encoding for $i$-th step, and $D$ indicates the hidden size. The positional encoding is defined as follows:
@@
The positional encoding helps GATSM effectively capture temporal patterns. While learnable position embedding also works in GATSM, we recommend positional encoding because position embedding requires knowledge of the maximum number of time steps, which is often unknown in real-world settings. After computing $\textbf{v}_{i}$, we calculate the attention scores as follows:
@@
where $k$ is attention head index, $\sigma\left(\cdot\right)$ is an activation function, $\textbf{w}_{k}^{attn}\in\mathbb{R}^{2D}$, and $m_{i,j}\in\mathbb{R}$ is the mask value used to block future information. The time mask is defined as follows:
@@
Inference: The prediction of GATSM is produced by combining the transformed features from time-sharing NBM with the attention scores from masked MHA.
@@
where $K$ is the number of attention heads, $\textbf{a}_{k,t}=\left[a_{k,i,1},a_{k,i,2},\cdots,a_{k,i,T}\right]\in\mathbb{R}^{T}$ is the attention map in Equation 8, $\tilde{\textbf{X}}=\left[\tilde{\textbf{x}}_{1},\tilde{\textbf{x}}_{2},\cdots,\tilde{\textbf{x}}_{T}\right]\in\mathbb{R}^{T\times M}$ is the transformed features in Equation 4, and $\textbf{w}_{k}^{out}\in\mathbb{R}^{M}$ is the learnable output weight.
Interpretability: We can rewrite Equation 10 as the following scalar form:
$\begin{split}\sum_{k=1}^{K}\textbf{a}_{k,t}^{\intercal}\tilde{\textbf{X}}\textbf{w}_{k}^{out}&=\sum_{u=1}^{t}\sum_{m=1}^{M}\sum_{k=1}^{K}\sum_{b=1}^{B}a_{k,t,u}h_{b}\left(x_{u,m}\right)w_{m,b}^{nbm}w_{k,m}^{out}\\ &=\sum_{u=1}^{t}\sum_{m=1}^{M}f_{u,m}\left(x_{u,m},\textbf{X}_{:t}\right)\end{split}$ (11)
Equation 11 shows that GATSM satisfying Definition 3.1. We can derive three types of interpretations from GATSM: 1) $a_{k,t,u}$ indicates the importance of time step $u$ at time step $t$, 2) $h_{b}\left(x_{u,m}\right)w_{m,b}^{nbm}w_{k,m}^{out}$ represents the time-independent contribution of feature $m$, and 3) $a_{k,t,u}h_{b}\left(x_{u,m}\right)w_{m,b}^{nbm}w_{k,m}^{out}$ represents the time-dependent contribution of feature $m$ at time step $t$.

### v3

@@ Abstract
Transparent models, which provide inherently interpretable predictions, are receiving significant attention in high-stakes domains. However, despite much real-world data being collected as time series, there is a lack of studies on transparent time series models. To address this gap, we propose a novel transparent neural network model for time series called Generalized Additive Time Series Model (GATSM). GATSM consists of two parts: 1) independent feature networks to learn feature representations, and 2) a transparent temporal module to learn temporal patterns across different time steps using the feature representations. This structure allows GATSM to effectively capture temporal patterns and handle varying-length time series while preserving transparency. Empirical experiments show that GATSM significantly outperforms existing generalized additive models and achieves comparable performance to black-box time series models, such as recurrent neural networks and Transformer. In addition, we demonstrate that GATSM finds interesting patterns in time series.
Codehttps://github.com/gim4855744/GATSM
Extended versionhttps://arxiv.org/abs/2410.10535
@@ 1 Introduction
Artificial neural networks excel at learning complex representations and demonstrate remarkable predictive performance across various fields. However, their complexity makes interpreting the decision-making processes of neural network models challenging. Consequently, researchers have widely studied post-hoc explainable artificial intelligence (XAI) methods, which explain the predictions of trained black-box models, in recent years (ribeiro_why_2016; lundberg_unified_2017; selvaraju_grad-cam_2017; mothilal_explaining_2020). XAI methods are generally effective at providing humans with understandable explanations of model predictions. However, they may produce incorrect and unfaithful explanations of the underlying black-box model and cannot provide actual contributions of input features to model predictions (cynthia_rudin_please_2018; cynthia_rudin_stop_2019; rahnama_study_2019; liu_synthetic_2021; akhavan_rahnama_blame_2023; rahnama_can_2024). Therefore, their applicability to high-stakes domains – such as healthcare and fraud detection, where faithfulness to the underlying model and actual contributions of features are important – is limited.
@@
Due to these limitations, transparent (i.e., inherently interpretable) models are attracting attention as alternatives to XAI in high-stakes domains (rishabh_agarwal_neural_2021; chun-hao_chang_node-gam_2022; filip_radenovic_neural_2022). Modern transparent models typically adhere to the generalized additive model (GAM) framework (hastie_generalized_1986). A GAM consists of independent functions, each corresponding to an input feature, and makes predictions as a linear combination of these functions (e.g., the sum of all functions). Therefore, each function reflects the contribution of its respective feature. For this reason, interpreting GAMs is straightforward, making them widely used in various fields, such as healthcare (rich_caruana_intelligible_2015; chang_how_2021), survival analysis (utkin_survnam_2022), and model bias discovery (rishabh_agarwal_neural_2021; tan_distill-and-compare_2018; kim_higher-order_2022). However, despite much real-world data being collected as time series, research on GAMs for time series remains scarce. Consequently, the applicability of GAMs in real-world scenarios is still limited.
@@
To overcome this limitation, we propose a novel transparent model for multivariate time series called Generalized Additive Time Series Model (GATSM). GATSM consists of independent feature networks to learn feature representations and a transparent temporal module to learn temporal patterns. Since employing distinct networks across different time steps requires a massive amount of learnable parameters, the feature networks in GATSM share the weights across all time steps, while the temporal module learns temporal patterns. GATSM then generates final predictions by integrating the feature representations with the temporal information from the temporal module. This strategy allows GATSM to effectively capture temporal patterns and handle dynamic-length time series while preserving transparency. Additionally, this approach facilitates the separate extraction of time-independent feature contributions, the importance of individual time steps, and time-dependent feature contributions through the feature functions, temporal module, and final prediction. To demonstrate the effectiveness of GATSM, we conducted empirical experiments on various real-world and synthetic time series datasets. The experimental results show that GATSM significantly outperforms existing GAMs and achieves comparable performance to black-box time series models, such as recurrent neural networks and Transformer (ashish_vaswani_attention_2017). In addition, we provide visualizations of GATSM’s predictions to demonstrate that GATSM finds interesting patterns in time series.
A detailed discussion of related works is provided in Appendix C and we summarize the contributions of this paper as follows: 1) We propose a novel transparent neural network model for multivariate time series, called GATSM. To the best of our knowledge, GATSM is the first transparent model capable of capturing temporal patterns while preserving transparency. 2) We conduct extensive experiments to validate the performance of GATSM. The experimental results demonstrate that GATSM significantly outperforms existing transparent models and achieves performance comparable to black-box models like Transformer. 3) We provide visual interpretations of GATSM’s predictions. These visualizations illustrate that GATSM successfully captures complex temporal patterns and derive multiple forms of interpretations, including time-step importance, global feature contributions, local time-independent feature contributions, and local time-dependent feature contributions. Such interpretations significantly enhance the understanding of both model behaviors and underlying dataset characteristics. 4) We thoroughly discuss the potential applications and limitations of GATSM, identifying directions for future works.
2 Problem Statement
@@
The simple linear model is designed to fit the conditional expectation $g\left(\mathbb{E}\left(y\mid\textbf{x}\right)\right)=\sum_{i=1}^{M}x_{i}w_{i}$, where $g(\cdot)$ is a link function, $M$ indicates the number of input features, $y$ is the target value for the given input features $\textbf{x}\in\mathbb{R}^{M}$, and $w_{i}\in\mathbb{R}$ is the learnable weight for $x_{i}$. This model captures only linear relationships between the target $y$ and the inputs x. To address this limitation, GAM (hastie_generalized_1986) extends the simple linear model to the generalized form as follows:
@@
where each $f_{i}(\cdot)$ is a function that models the effect of a single feature, referred as a feature function. Typically, $f_{i}\left(\cdot\right)$ becomes a non-linear function such as a decision tree or neural network to capture non-linear relationships. However, this form of GAM cannot handle time series data. One straightforward method to extend GAM to time series, adopted in NATM (jo_neural_2023), is applying distinct feature functions to each time step and summing them to produce predictions:
where $\textbf{X}\in\mathbb{R}^{T\times M}$ is a time series with $T$ time steps and $M$ features, and $t$ is the current time step. This method can handle time series data as input but fails to capture temporal patterns because $f_{i,j}\left(\cdot\right)$ still does not interact with previous time steps. To overcome this problem, we suggest a new form of GAM for time series.
Definition 3.1 GAMs for time series, which capture temporal patterns hold the following form:
In Equation (3), $f_{i,j}\left(\cdot,\cdot\right)$ can capture interactions between current and previous time steps. Therefore, GAMs adhering to Definition 3.1 can capture temporal patterns. However, implementing such a model while maintaining transparency poses challenges. In the following section, we will describe our approach to implementing a GAM that satisfies Definition 3.1. To the best of our knowledge, no existing literature addresses Definition 3.1. Table 1 shows the advantages of our GATSM compared to existing GAMs.
Time series Temporal pattern Dynamic length
Tabular GAMs
@@
Table 1: Advantages of GATSM. Unlike existing tabular GAMs that cannot handle time series data, and NATM which only supports fixed-length time series without capturing temporal dynamics, GATSM effectively handles varying-length time series and captures complex temporal patterns.
Figure 1: Architecture of GATSM. GATSM comprises two components: Time-Sharing NBM and Masked MHA. Time-Sharing NBM learns nonlinear representations of input features without considering temporal patterns, using parameters shared across time steps. Masked MHA then captures temporal patterns based on these feature representations.
3 Generalized Additive Time Series Model
Figure 1 shows the overall architecture of GATSM. Our model has two modules: 1) feature networks, called time-sharing neural basis model (NBM), for learning feature representations, and 2) masked multi-head attention (MHA) for learning temporal patterns.
Time-Sharing NBM
Assume a time series $\textbf{X}\in\mathbb{R}^{T\times M}$ with $T$ time steps and $M$ features. Applying existing GAMs defined in Equation 1 to this time series require $T\times M$ feature functions, which becomes problematic when dealing with large $T$ or $M$ due to increased model size. This limits the applicability of GAMs to real-world datasets. To overcome this problem, we extend NBM (filip_radenovic_neural_2022) to time series as:
@@ 3 Generalized Additive Time Series Model
We refer to this extended version of NBM as time-sharing NBM. Time-sharing NBM has $B$ basis functions, with each basis function $h_{k}(\cdot)$ taking a feature $x_{i,j}$ as input. The feature-specific weight $w_{j,k}^{nbm}$ then projects the basis to the transformed feature $\tilde{x}_{i,j}$. As Equation (4) depicts, the basis functions are shared across all features and time steps, drastically reducing the number of required feature functions from $T\times M$ to $B$. We use $B=100$ and implement $h_{k}\left(\cdot\right)$ using multi-layer perceptron (MLP).
Masked MHA
GATSM employs MHA to learn temporal patterns. Although the dot product attention (ashish_vaswani_attention_2017) is popular, the simple dot product attention has low expressive power (velickovic_graph_2018). Therefore, we adopt the 2-layer attention mechanism proposed by (velickovic_graph_2018) to GATSM. We first transform $\tilde{\textbf{x}}_{i}=\left[\tilde{x}_{i,1},\tilde{x}_{i,2},\cdots,\tilde{x}_{i,M}\right]\in\mathbb{R}^{M}$ produced by Equation (4) as follows:
@@
where $\textbf{Z}\in\mathbb{R}^{M\times D}$ is a learnable weight, $\textbf{pe}_{i}=\left[pe_{i,1},pe_{i,2},\cdots,pe_{i,D}\right]\in\mathbb{R}^{D}$ is the positional encoding for the $i$-th step, and $D$ indicates the hidden size. The positional encoding is defined as follows:
@@
The positional encoding helps GATSM effectively capture temporal patterns. While learnable position embedding also works in GATSM, we recommend the sinusoidal positional encoding because learnable position embedding requires prior knowledge of the maximum length of a time series, which is often unknown in real-world settings. For example, in real-world hospital settings, a patient’s length of stay continuously increases, and the discharge date is unpredictable. After computing $\textbf{v}_{i}$, we calculate the attention scores as follows:
@@
where $k$ is the index of the attention head, $\sigma\left(\cdot\right)$ is a non-linear activation function, $\textbf{w}_{k}^{attn}\in\mathbb{R}^{2D}$ is a learnable weight, and $m_{i,j}\in\mathbb{R}$ is the mask value used to block future information. The time mask is defined as follows:
@@
Inference
GATSM produces its prediction by combining the transformed features from the time-sharing NBM with the attention scores from the masked MHA.
@@
where $K$ is the number of attention heads, $\textbf{a}_{k,t}=\left[a_{k,i,1},a_{k,i,2},\cdots,a_{k,i,T}\right]\in\mathbb{R}^{T}$ is the attention map defined in Equation (8), $\tilde{\textbf{X}}=\left[\tilde{\textbf{x}}_{1},\tilde{\textbf{x}}_{2},\cdots,\tilde{\textbf{x}}_{T}\right]\in\mathbb{R}^{T\times M}$ are the transformed features defined in Equation (4), and $\textbf{w}_{k}^{out}\in\mathbb{R}^{M}$ is the learnable output weight.
Interpretability
We can rewrite Equation (10) as the following scalar form:
$\begin{split}\sum_{k=1}^{K}&\textbf{a}_{k,t}^{\intercal}\tilde{\textbf{X}}\textbf{w}_{k}^{out}\\ &=\sum_{u=1}^{t}\sum_{m=1}^{M}\sum_{k=1}^{K}\sum_{b=1}^{B}a_{k,t,u}h_{b}\left(x_{u,m}\right)w_{m,b}^{nbm}w_{k,m}^{out}\\ &=\sum_{u=1}^{t}\sum_{m=1}^{M}f_{u,m}\left(x_{u,m},\textbf{X}_{:t}\right)\end{split}$ (11)
Equation (11) shows that GATSM satisfies Definition 3.1 by encapsulating the attention scores (i.e., $\alpha_{k,t,u}$) and time-sharing NBM (i.e., $h_{b}$) into a function $f_{u,m}$. We can derive three types of interpretations from GATSM: 1) $a_{k,t,u}$ indicates the importance of time step $u$ at time step $t$, 2) $h_{b}\left(x_{u,m}\right)w_{m,b}^{nbm}w_{k,m}^{out}$ represents the time-independent contribution of feature $m$, and 3) $a_{k,t,u}h_{b}\left(x_{u,m}\right)w_{m,b}^{nbm}w_{k,m}^{out}$ represents the time-dependent contribution of feature $m$ at time step $t$.